OpenStudy (yessenia27):

Raul calculated that he would spend $125 on school supplies this year. He actually spent $87.50 on school supplies. What is Raul’s percent of error?

OpenStudy (xapproachesinfinity):

percentage error is calculated like this original value-true value divided by the original value times 100

OpenStudy (yessenia27):

What ?

OpenStudy (xapproachesinfinity):

this is what you have to do \[\frac{125-87.5}{125}\times 100 \]

OpenStudy (xapproachesinfinity):

actually a mistake it is \[\frac{125-87.5}{87.5}\times 100\]

OpenStudy (xapproachesinfinity):

42% is the error